OpenStudy (anonymous):

Find the antiderivative. f(x) = x^2

OpenStudy (lgbasallote):

use power rule

OpenStudy (lgbasallote):

\[\large \int x^n dx = \frac{x^{n+1}}{n+1}\] do you get that? or do you need more help?

OpenStudy (anonymous):

more help please

OpenStudy (anonymous):

If you compare your equation to his, what is n?

OpenStudy (lgbasallote):

hint: n is the exponent of x in my example

OpenStudy (anonymous):

hmm sorry im still confused

OpenStudy (lgbasallote):

what is the exponent in this problem?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

2x

OpenStudy (lgbasallote):

2 is the exponent right so what did i do with the exponent in \[x^n dx = \frac{x^{n+1}}{n+1}\]

OpenStudy (anonymous):

@student92 anti-derivative, not the derivative. Also, please don't just give out the answer without anything else as that's against the code of conduct: http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

put it where n is

OpenStudy (lgbasallote):

okay here's a more visual hint \[\large \int x^3 dx = \frac{x^{3+1}}{3+1} = \frac{x^4}{4}\]

OpenStudy (lgbasallote):

do you see it?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but thats not the antiderivative right?

OpenStudy (anonymous):

It's the antiderivative to the example he posted. To prove it you can take the derivative of (X^4}/4 and see if you get X^3 as the result.

OpenStudy (anonymous):

oh got it

OpenStudy (lgbasallote):

so do you now know what is \[\int x^2 dx?\]

OpenStudy (anonymous):

no...

OpenStudy (lgbasallote):

okay another hint \[\huge \int x^4 dx \implies \frac{x^{4+1}}{4+1} \implies \frac{x^5}{5}\]

OpenStudy (anonymous):

so thats the antiderivative

OpenStudy (anonymous):

?

OpenStudy (anonymous):

It's the anti-derivative of his example. Do you know what an anti-derivative is?

OpenStudy (anonymous):

yes, inverse of the derivative

OpenStudy (lgbasallote):

it just follows the rule \[\int x^n dx = \frac{x^{n+1}}{n+1}\] n is the exponent

OpenStudy (lgbasallote):

here's another one... \[\large \int xdx \implies \int x^1 dx \implies \frac{x^{1+1}}{1+1} \implies \frac{x^2}{2}\]

OpenStudy (lgbasallote):

do you see what is \[\int x^2 dx\] yet?

OpenStudy (anonymous):

yes i believe so

OpenStudy (zzr0ck3r):

int(x^3) = (1/4)x^4 int(x^5) = (1/6)x^6 int(x^32) = (1/33)x^33 int(x^2) = ?

OpenStudy (anonymous):

(1/3)x^3?

OpenStudy (zzr0ck3r):

:)

OpenStudy (anonymous):

awesome thanks!!!!

OpenStudy (zzr0ck3r):

fo sure

OpenStudy (anonymous):

Now try this one. What function has antiderivative = derivative of the function. If you can work that out. You'll understand how awesome e is.

OpenStudy (zzr0ck3r):

now take the derivative of that just for fun.

OpenStudy (zzr0ck3r):

prooving that said function is the ONLY antiderivative = derivative is sort of hard at this point im thinking